Circular No. 5740

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993J IN NGC 3031 A. V. Filippenko and T. Matheson, University of California at Berke-

ley, report that careful examination of the fully reduced spectrum obtained on Mar. 30.3 UT (cf. IAUC 5731) reveals many broad undulations with peak-to-valley amplitudes of 2-8 percent. The peak of the most prominent bump is near 536 nm; others are present at 450, 561, 645 nm and elsewhere. Adjacent to the 645-nm emission, the amplitude of which is only 3 percent, is a broad absorption trough of comparable strength with a minimum near 627 nm; overall, this feature resembles a weak P-Cyg profile. If this is identified with Hα, the expansion velocity derived from the minimum in the absorption trough is 13 000 km/s, and the emission peak is blueshifted by 5000 km/s because we are viewing predominantly the near side of opticallythick ejecta. Several weak, unresolved (at 0.5-nm resolution) emission lines are also visible in the spectrum. These include Hα at an observed (heliocentric) wavelength of 656.00 nm (equivalent width 0.014 nm), He II at 468.36 nm (EW 0.035 nm; possibly resolved), and a line at 637.14 nm (EW 0.034 nm) that might be [Fe X] 637.4 nm. High-excitation emission lines could be produced by circumstellar gas photoionized by ultraviolet radiation from the hot stellar surface after breakout of the supernova shock wave. Unresolved interstellar absorption lines are visible at 373.45 nm (EW 0.028 nm; possibly Al II 373.46), 393.34 (EW 0.066 nm; Ca II K) and 396.82 (EW 0.038 nm; Ca II H). The interstellar Na I D lines are blended, perhaps due to the presence of multiple components and have total EW 0.16 nm. There is also an unidentified, resolved absorption line at 615.2 nm (EW 0.05 nm; FWHM 1.7 nm) and several possibly resolved absorption lines blueward of the narrow Hα emission. Given the weakness of the broad features, it is possible that at least some of them are produced by subtle calibration errors. On the other hand, R. P. Kirshner and B. P. Schmidt, Center for Astrophysics, confirm those at 450 and 540 nm, as well as the narrow He II, Hα and 637-nm emission lines, in partially calibrated spectra (ranges 380-570 nm, 620-720 nm) obtained by N. Caldwell with the Multiple-Mirror Telescope on Mar. 31.12 UT.”

The object’s outburst seems to have occurred between Mar. 27.91 UT, when J.-C. Merlin, Le Creusot, France, failed to record the supernova to m_{v} = 16.0 on an Ektachrome 400 exposure, and Mar. 28.30 UT, when A. Neely, Silver City, NM, found it at mag 13.8 on an unfiltered CCD image.

Visual magnitude estimates: Apr. 1.08 UT, 10.7 (P. Schmeer, Bischmisheim, Germany); 1.81, 10.5 (T. Vanmunster, Landen, Belgium).

1993 April 2 Brian G. Marsden